OpenStudy (anonymous):

Y is between X and N. XN=10 and YX=1 2/3. Find YN.

OpenStudy (anonymous):

@Learner11

OpenStudy (anonymous):

Here XN=10 and YX=1 2/3 Then .  So,YN=XN-YX=10-1 2/3 = 8.33